# LW-FedSSL: Resource-Efficient Layer-wise Federated Self-Supervised Learning

### v2

@@ VI Conclusion
Edge devices in distributed environments often struggle to meet the computation and communication requirements of FedSSL. In response to these challenges, we propose a layer-wise training approach called LW-FedSSL, enabling FL clients to reduce their resource requirements. Our experiments demonstrate that LW-FedSSL has a $\times3$ lower memory requirement, 2.1×2.1$\times$2.1 × lower total FLOPs consumption, and 3.2×3.2$\times$3.2 × cheaper total transmission cost (download + upload), while achieving comparable performance to its end-to-end training counterpart. Furthermore, we explore a progressive training approach called Prog-FedSSL, which remains more resource-efficient than end-to-end training, with significantly improved performance.
@@ Appendix A Additional Details
Figure 15: Comparison between FedMoCo (left), LW-FedMoCo (center), and Prog-FedSSL (right) at stage $s$.
@@
A-A Self-supervised Learning
Algorithm 3 describes the detailed training procedure of MoCo v3 [11].
A-B Data Distribution
Fig. 16 illustrates the data distribution among clients for different $\beta$ values used in Section V-F. Lower $\beta$ values correspond to increasingly heterogeneous data distributions among clients.
TABLE V: Linear evaluation results of LW-FedSSL with different $\alpha$ values.
@@ A-B Data Distribution
TABLE VI: Linear evaluation results with and without weight transfer (WT).
@@
A-C Details on FLOPs Calculation
Here, we discuss the details on FLOPs calculation used in Section V-B. In layer-wise training, calculating FLOPs for inactive (frozen) layers involves simply considering FLOPs associated with the forward pass. However, for active layers, the calculation must account for both the forward and backward passes. While numerous works have addressed FLOPs calculation for the forward pass (inference), limited studies are available on the practice of FLOPs calculation for the backward pass. Notably, existing works related to layer-wise training [18, 19] also do not provide specific details on how FLOPs are computed. Some studies [40, 41, 42, 43] suggest that the number of operations in a backward pass of a neural network is often twice that of a forward pass. Following these studies, we adopt a backward-forward ratio of 2:1 to calculate the total FLOPs for active layers. We use the FlopCountAnalysis tool in the fvcore [44] library for calculating the FLOPs of the forward pass. We only consider a single input sample for FLOPs calculation.
Appendix B Additional Experiments
TABLE VII: Projection head $H$.
@@ Appendix B Additional Experiments
TABLE VIII: Prediction head $P$.
@@
B-A Weight of Representation Alignment
In LW-FedSSL, the weight term $\alpha$ controls the regularization strength of the representation alignment mechanism in the local training process. We explore different $\alpha$ values in Table V. Based on the results, we set the default value of $\alpha$ to 0.01.
B-B Weight Transfer
For LW-FedMoCo, LW-FedSSL, and Prog-FedSSL, we transfer the weights of the prior layer, $L_{s-1}$, to the subsequent layer, $L_{s}$, at the beginning of a new stage $s$. Here, we investigate the impact of weight transfer on the performance of these approaches. Table VI demonstrates that weight transfer can improve the linear evaluation performance of the baseline LW-FedMoCo across different downstream datasets. For LW-FedSSL, weight transfer improves the performance on STL-10 and CIFAR-100 but decreases on CIFAR-10 and Tiny ImageNet datasets. Similarly, in the case of Prog-FedSSL, weight transfer leads to improved performance on STL-10 but a slight decrease on Tiny ImageNet, while maintaining comparable performance on other datasets. Notably, weight transfer consistently improves performance across all approaches on the STL-10 dataset, a case where the downstream data originates from a similar distribution as the unlabeled pre-training data within clients. This highlights a potential utility of weight transfer in the context of layer-wise training, where downstream data and pre-training data originate from a similar distribution (e.g., semi-supervised scenarios where only a portion of the data is labeled).
B-C Hidden Layer Dimension in MLP Heads
Following [11], we set the dimension of hidden layers in the projection head $H$ and the prediction head $P$ to 4096 by default. Tables VII and VIII describe the detailed structures of $H$ and $P$, respectively. We further explore various dimensions, specifically 128, 256, 512, and 1024. The results in Fig. 17 indicate that Prog-FedSSL is relatively sensitive to the hidden layer dimension of the MLP heads.

### v3

@@ 6 Conclusion
Edge devices in distributed environments often struggle to meet the computation and communication requirements of FedSSL. In response to these challenges, we propose a layer-wise training approach named LW-FedSSL, which allows FL clients to perform self-supervised learning with improved resource efficiency. By leveraging server-side calibration and representation alignment mechanisms, LW-FedSSL achieves comparable performance to end-to-end training while significantly reducing resource demands. Specifically, it requires $\times3$ less memory, consumes 2.1×2.1$\times$2.1 × fewer FLOPs, and reduces total transmission costs (download and upload) by 3.2×3.2$\times$3.2 ×. Furthermore, we explore a progressive training approach called Prog-FedSSL. Although it offers lower resource efficiency than LW-FedSSL, it remains more efficient than end-to-end training, while also improving performance.
@@ Appendix A Additional Details
A.1 Details on FLOPs Calculation
Here, we discuss the details on FLOPs calculation used in Section 5.2. In layer-wise training, calculating FLOPs for inactive (frozen) layers involves simply considering FLOPs associated with the forward pass. However, for active layers, the calculation must account for both the forward and backward passes. While numerous works have addressed FLOPs calculation for the forward pass (inference), limited studies are available on the practice of FLOPs calculation for the backward pass. Notably, existing works related to layer-wise training [24, 25] also do not provide specific details on how FLOPs are computed. Some studies [44, 45, 46, 47] suggest that the number of operations in a backward pass of a neural network is often twice that of a forward pass. Following these studies, we adopt a backward-forward ratio of 2:1 to calculate the total FLOPs for active layers. We use the FlopCountAnalysis tool in the fvcore [48] library for calculating the FLOPs of the forward pass. We only consider a single input sample for FLOPs calculation.
A.2 Data Distribution
Fig. 16 illustrates the data distribution among clients for different $\beta$ values used in Section 5.6. Lower $\beta$ values correspond to increasingly heterogeneous data distributions among clients.
A.3 Self-supervised Learning
Algorithm 3 describes the detailed training procedure of MoCo v3 [11].
@@ A.3 Self-supervised Learning
Figure 15: Comparison between FedMoCo, FedMoCo-LW, and Prog-FedSSL at stage $s$.
@@
Appendix B Additional Experiments
B.1 Weight of Representation Alignment
In LW-FedSSL, the weight term $\alpha$ controls the regularization strength of the representation alignment mechanism in the local training process. We explore different $\alpha$ values in Table 5. Based on the results, we set the default value of $\alpha$ to 0.01.
Table 5: Linear evaluation results of LW-FedSSL with different $\alpha$ values.
@@ B.1 Weight of Representation Alignment
Table 6: Linear evaluation results with and without weight transfer (WT).
@@
B.2 Weight Transfer
For FedMoCo-LW, LW-FedSSL, and Prog-FedSSL, we transfer the weights of the prior layer, $L_{s-1}$, to the subsequent layer, $L_{s}$, at the beginning of a new stage $s$. Here, we investigate the impact of weight transfer on the performance of these approaches. Table 6 demonstrates that weight transfer can improve the linear evaluation performance of the baseline FedMoCo-LW across different downstream datasets. For LW-FedSSL, weight transfer improves the performance on STL-10 and CIFAR-100 but decreases on CIFAR-10 and Tiny ImageNet datasets. Similarly, in the case of Prog-FedSSL, weight transfer leads to improved performance on STL-10 but a slight decrease on Tiny ImageNet, while maintaining comparable performance on other datasets. Notably, weight transfer consistently improves performance across all approaches on the STL-10 dataset, a case where the downstream data originates from a similar distribution as the unlabeled pre-training data within clients. This highlights a potential utility of weight transfer in the context of layer-wise training, where downstream data and pre-training data originate from a similar distribution (e.g., semi-supervised scenarios where only a portion of the data is labeled).
B.3 Hidden Layer Dimension in MLP Heads
Following [11], we set the dimension of hidden layers in the projection head $H$ and the prediction head $P$ to 4096 by default. Tables 7 and 8 describe the detailed structures of $H$ and $P$, respectively. We further explore various hidden dimensions, specifically 128, 256, 512, and 1024. The results in Fig. 17 indicate that Prog-FedSSL is relatively sensitive to the hidden layer dimension of the MLP heads.
Table 7: Projection head $H$.
@@ B.3 Hidden Layer Dimension in MLP Heads
Table 8: Prediction head $P$.